# Stellar Superradiance and Low-Energy Absorption in Dense Nuclear Media

### v1

@@ 1 Introduction
Hypothetical bosons such as axions and dark photons are well-motivated extensions beyond the Standard Model (SM) of particle physics. The QCD axion provides an elegant solution to the strong CP problem [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], while broader families of axion-like particles and dark photons arise generically in fundamental theories with extra dimensions [11, 12, 13, 14]. A wide range of astrophysical observations place strong constraints on such bosons. In particular, stellar-cooling (SC) arguments exclude large regions of parameter space for boson masses below the eV scale by requiring that new dissipative channels do not exceed energy loss from Standard Model emission processes [15, 16, 17]. These bounds come from a diverse of environments, including neutron star (NS) and supernovae cooling constraints on couplings to nucleons [18, 19, 20], as well as white dwarf [21] and red giant [22] limits on interactions with electrons.
@@ 3.2.2 Long-Wavelength Suppression in Dense Nuclear Media
The suppression factor, $F_{\rm sup}$, arises from the fact that neutrons in a dense medium undergo many scatterings during a single oscillation of the bosonic bound state. In this regime, the boson cannot resolve individual neutron–neutron interactions; instead, absorption and emission couple to the collective response of the medium. Quantitatively, repeated NN scatterings broaden the neutron propagator by introducing a collision width $\Gamma_{\rm col}$ [57, 58, 59, 60, 61, 15]. This modifies the in-medium amplitude in close analogy with the Landau–Pomeranchuk–Migdal effect [82, 83], yielding the suppression factor
$F_{\rm sup}=\frac{\omega_{b}^{2}}{\omega_{b}^{2}+\Gamma_{\rm col}^{2}}.$ (3.11)
The factor of $\omega_{b}$ in the numerator reflects the gradient or time derivative acting on the bosonic field, while the $\omega_{b}$ appearing in the denominator arises from the free kinetic part of the neutron propagator connected to the boson line in Fig. 1. When $\omega_{b}\ll\Gamma_{\rm col}$, the boson effectively probes only an averaged, rapidly fluctuating neutron medium, leading to strong suppression: $F_{\rm sup}\simeq\omega_{b}^{2}/\Gamma_{\rm col}^{2}$.
@@ 4 Discussion
Although our analysis has focused on axion and dark-photon couplings to neutrons, similar considerations apply to a broader class of interactions, including quadratic couplings [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121] and axion–EDM couplings [122, 123]. These scenarios would likewise require a careful treatment of dense-matter effects and long-wavelength behavior.
@@ Acknowledgments
$\mathcal{L}_{\pi N}=\frac{g_{\pi N}}{m_{\pi}}(\partial_{\mu}\pi^{0})\,\bar{\psi}\gamma^{\mu}\gamma^{5}\psi,$ (S1)
@@
$\left\{\mathcal{M}_{k}^{2}\right\}\equiv\frac{|\vec{k}|^{4}}{(|\vec{k}|^{2}+m_{\pi}^{2})^{2}}\,,\\ \left\{\mathcal{M}_{l}^{2}\right\}\equiv\frac{|\vec{l}|^{4}}{(|\vec{l}|^{2}+m_{\pi}^{2})^{2}}\,,\\ \left\{\mathcal{M}_{kl}^{2}\right\}\equiv\frac{|\vec{k}|^{2}|\vec{l}|^{2}}{(|\vec{k}|^{2}+m_{\pi}^{2})(|\vec{l}|^{2}+m_{\pi}^{2})}\,.$ (S2)
1 Stellar Cooling
@@
$\sum_{s}|\mathcal{M}^{a}_{\mathrm{SC}}|^{2}$ $=\frac{256}{3}\,\frac{g_{\pi N}^{4}\,m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),$ (S3)
$\sum_{s}|\mathcal{M}^{\mathrm{EDM/MDM}}_{\mathrm{SC}}|^{2}$ $=\frac{512}{3}\,\frac{g_{\pi N}^{4}\,m_{\psi}^{2}\,g_{\mathrm{EDM/MDM}}^{\,2}}{m_{\pi}^{4}}\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),$ (S4)
@@
$Q=\int{\rm d}\Pi_{b}\,\omega_{b}\!\left[\prod_{i=1}^{4}\!\int\frac{{\rm d}^{3}\vec{p}_{i}}{2\omega_{i}(2\pi)^{3}}\right]f_{1}f_{2}(1-f_{3})(1-f_{4})(2\pi)^{4}\delta^{4}(p_{1}+p_{2}-p_{3}-p_{4}-p_{b})\,\frac{\sum_{s}|\mathcal{M}_{\mathrm{SC}}|^{2}}{4},$ (S5)
@@
${\rm d}^{3}\vec{p}_{i}=|\vec{p}_{i}|^{2}{\rm d}|\vec{p}_{i}|{\rm d}\Omega_{i}\simeq p_{F}m_{\psi}{\rm d}\omega_{i}{\rm d}\Omega_{i}=(m_{\psi}p_{F}T)\,{\rm d}y_{i}{\rm d}\Omega_{i},$ (S6)
@@
$\begin{split}Q_{a}&=\frac{p_{F}^{4}}{2^{16}\pi^{10}}\!\left(\frac{256}{3}\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\right)A^{\mathrm{SC}}B^{\mathrm{SC}},\\[3.0pt] Q_{\mathrm{EDM/MDM}}&=\frac{p_{F}^{4}}{2^{16}\pi^{10}}\!\left(\frac{512}{3}\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{\mathrm{EDM/MDM}}^{\,2}}{m_{\pi}^{4}}\right)A^{\mathrm{SC}}B^{\mathrm{SC}},\end{split}$ (S7)
@@
$\begin{split}A^{\mathrm{SC}}&=\!\left[\prod_{i=1}^{4}\!\int{\rm d}\Omega_{i}\right]\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\!\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),\\[5.0pt] B^{\mathrm{SC}}&=T^{4}\!\!\int_{0}^{\infty}\!\!{\rm d}\omega_{b}\,\omega_{b}^{2}\!\left[\prod_{i=1}^{4}\!\int_{-\mu_{i}/T}^{\infty}\!\!{\rm d}y_{i}\right]\!\delta(\omega_{1}+\omega_{2}-\omega_{3}-\omega_{4}-\omega_{b})f_{1}f_{2}(1-f_{3})(1-f_{4}),\end{split}$ (S8)
@@
$B^{\mathrm{SC}}$ $\approx T^{6}\!\!\int_{0}^{\infty}\!{\rm d}y_{b}\,y_{b}^{2}\!\left[\prod_{i=1}^{4}\!\int_{-\infty}^{\infty}\!{\rm d}y_{i}\,f_{i}\right]\delta\!\left(\sum_{i}y_{i}-y_{b}\right)=\frac{T^{6}}{6}\!\int_{0}^{\infty}\!{\rm d}y_{b}\,\frac{y_{b}^{3}(y_{b}^{2}+4\pi^{2})}{e^{y_{b}}-1}=\frac{62\pi^{6}T^{6}}{945}.$ (S9)
@@
$F(\xi)$ $=4-\frac{1}{1+\xi^{2}}+5\xi\arctan(\xi)+\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}}\arctan\!\left(\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}}\right)+\frac{\pi}{2}\!\left(\frac{\xi}{\sqrt{2\xi^{2}+1}}-5\right)\xi.$ (S11)
@@ 2 Stellar Superradiance
$\begin{split}&\sum_{s}\left|\mathcal{M}^{a}_{\rm SS}\right|^{2}=\Bigg[\left\{\sum_{i=1}^{4}|\vec{v}_{i}|^{2}-4(\hat{l}\cdot\vec{v}_{1}+\hat{l}\cdot\vec{v}_{3})(\hat{l}\cdot\vec{v}_{2}+\hat{l}\cdot\vec{v}_{4})-4(\hat{k}\cdot\vec{v}_{1}+\hat{k}\cdot\vec{v}_{3})(\hat{k}\cdot\vec{v}_{2}+\hat{k}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{kl}^{2}\right\}\\ &+\sum_{i=1}^{4}|\vec{v}_{i}|^{2}\left(\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}\right)+\left\{2(\hat{k}\cdot\vec{v}_{1})(\hat{k}\cdot\vec{v}_{3})+2(\hat{k}\cdot\vec{v}_{2})(\hat{k}\cdot\vec{v}_{4})-(\vec{v}_{1}\cdot\vec{v}_{3})-(\vec{v}_{2}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{k}^{2}\right\}\\ &+\left\{2(\hat{l}\cdot\vec{v}_{2})(\hat{l}\cdot\vec{v}_{3})+2(\hat{l}\cdot\vec{v}_{1})(\hat{l}\cdot\vec{v}_{4})-(\vec{v}_{2}\cdot\vec{v}_{3})-(\vec{v}_{1}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{l}^{2}\right\}\Bigg]\,\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.\end{split}$ (S12)
@@
$\braket{\vec{k}\cdot\vec{q}_{b}}=\braket{\vec{l}\cdot\vec{q}_{b}}=\braket{\vec{p}_{1}\cdot\vec{q}_{b}}=0,\quad$ (S13)
@@
$\left\langle\sum_{\text{s}}\left|\mathcal{M}^{a}_{\rm SS}\right|^{2}\right\rangle=\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right]\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.$ (S14)
@@
$\varepsilon_{+1}^{\mu}(q)\varepsilon_{+1}^{*\nu}(q)\approx\frac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&i&0\\ 0&-i&1&0\\ 0&0&0&0\end{pmatrix}$, (S15)
@@
$\left\langle k_{x}\,p_{1x}\right\rangle=-\frac{|\vec{k}|^{2}}{6},$ (S16)
@@
$\left\langle\sum_{s}\big|\mathcal{M}^{\rm MDM}_{\rm SS}\big|^{2}\right\rangle=$ $\;\Bigg\{\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}\bigg[\frac{4}{3}\{\mathcal{M}_{k}^{2}\}+\frac{4}{3}\{\mathcal{M}_{l}^{2}\}+\frac{8}{9}\{\mathcal{M}_{kl}^{2}\}\bigg]+\frac{2}{3}\frac{p_{F}^{2}}{m_{\psi}^{2}}\big[\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}\big]$ (S17)
@@
$\left\langle\sum_{s}|\mathcal{M}^{\mathrm{EDM}}_{\rm SS}|^{2}\right\rangle=\frac{256}{3}\left(\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right)\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm EDM}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.$ (S18)
2.1 Phase Space Integral

### v2

@@ 1 Introduction
Hypothetical bosons such as axions and dark photons are well-motivated extensions beyond the Standard Model (SM) of particle physics. The QCD axion provides an elegant solution to the strong CP problem [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], while broader families of axion-like particles and dark photons arise generically in fundamental theories with extra dimensions [11, 12, 13, 14]. A wide range of astrophysical observations place strong constraints on such bosons. In particular, stellar-cooling (SC) arguments exclude large regions of parameter space for boson masses below the eV scale by requiring that new dissipative channels do not exceed energy loss from Standard Model emission processes [15, 16, 17]. These bounds come from a diverse of environments, including neutron star (NS) and supernovae cooling constraints on couplings to nucleons [18, 19, 20, 21, 22], as well as white dwarf [23] and red giant [24] limits on interactions with electrons. Axions emitted from these astrophysical sources can be further converted into gamma-ray photons in astrophysical magnetic fields through axion-photon coupling. This process leads to significant constraints on the axion parameter space using gamma-ray telescopes [25, 26, 27, 28, 29, 30].
@@ 3.2.2 Long-Wavelength Suppression in Dense Nuclear Media
The suppression factor $F_{\rm sup}$ originates from the fact that, in dense nuclear matter, neutrons undergo many NN scatterings during a single oscillation period of the bosonic bound state. When the boson frequency $\omega_{b}$ is small compared to the typical NN collision rate, the boson cannot resolve individual scattering events; instead, absorption and emission probe only the collective, rapidly fluctuating neutron medium. This physics is closely analogous to the Landau–Pomeranchuk–Migdal effect in bremsstrahlung processes [90, 91].
A quantitative understanding follows directly from Fig. 1 and the standard structure of NN bremsstrahlung in dense matter. Radiation emission or absorption factorizes into an NN scattering kernel and a radiation vertex attached to one of the nucleon legs. In the absence of medium effects, the intermediate nucleon propagator between the NN collision and the radiation vertex is nearly on shell, producing the familiar soft-radiation enhancement $\propto 1/\omega_{b}$, where $\omega_{b}$ is the energy transfer carried by the boson. Physically, this corresponds to a formation time $\sim 1/\omega_{b}$ during which the nucleon resides in a virtual off-shell state.
In a dense medium, however, additional NN scatterings typically occur within this formation time. As shown in classic studies of NN bremsstrahlung [65, 66, 67, 68, 69, 15], these multiple scatterings are incorporated by dressing the intermediate neutron propagator with the imaginary part of its in-medium self-energy, introducing a collision width $\Gamma_{\rm col}$. At the level of the propagator, this amounts to the replacement
$\frac{1}{\omega_{b}}\;\longrightarrow\;\frac{1}{\omega_{b}+i\,\Gamma_{\rm col}}.$ (3.11)
For bosonic absorption or emission, the radiation vertex involves a gradient or time derivative acting on the bosonic field, supplying an explicit factor of $\omega_{b}$ in the numerator of the amplitude. Combining this with the damped propagator, the squared in-medium amplitude acquires the universal suppression factor
$F_{\rm sup}=\left|\frac{\omega_{b}}{\omega_{b}+i\Gamma_{\rm col}}\right|^{2}=\frac{\omega_{b}^{2}}{\omega_{b}^{2}+\Gamma_{\rm col}^{2}}.$ (3.12)
In the regime $\omega_{b}\ll\Gamma_{\rm col}$, many NN collisions occur within a single boson oscillation period, the intermediate neutron state loses phase coherence, leading to strong suppression, $F_{\rm sup}\simeq\omega_{b}^{2}/\Gamma_{\rm col}^{2}$.
@@ 4 Discussion
Although our analysis has focused on axion and dark-photon couplings to neutrons, similar considerations apply to a broader class of interactions, including quadratic couplings [110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128] and axion–EDM couplings [129, 130]. These scenarios would likewise require a careful treatment of dense-matter effects and long-wavelength behavior [131].
@@ Acknowledgments
$\mathcal{L}_{\pi N}=\frac{g_{\pi N}}{m_{\pi}}(\partial_{\mu}\pi^{0})\,\bar{\psi}\gamma^{\mu}\gamma^{5}\psi,$ (.1)
@@
$\left\{\mathcal{M}_{k}^{2}\right\}\equiv\frac{|\vec{k}|^{4}}{(|\vec{k}|^{2}+m_{\pi}^{2})^{2}}\,,\\ \left\{\mathcal{M}_{l}^{2}\right\}\equiv\frac{|\vec{l}|^{4}}{(|\vec{l}|^{2}+m_{\pi}^{2})^{2}}\,,\\ \left\{\mathcal{M}_{kl}^{2}\right\}\equiv\frac{|\vec{k}|^{2}|\vec{l}|^{2}}{(|\vec{k}|^{2}+m_{\pi}^{2})(|\vec{l}|^{2}+m_{\pi}^{2})}\,.$ (.2)
Appendix A Stellar Cooling
@@
$\sum_{s}|\mathcal{M}^{a}_{\mathrm{SC}}|^{2}$ $=\frac{256}{3}\,\frac{g_{\pi N}^{4}\,m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),$ (A.1)
$\sum_{s}|\mathcal{M}^{\mathrm{EDM/MDM}}_{\mathrm{SC}}|^{2}$ $=\frac{512}{3}\,\frac{g_{\pi N}^{4}\,m_{\psi}^{2}\,g_{\mathrm{EDM/MDM}}^{\,2}}{m_{\pi}^{4}}\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),$ (A.2)
@@
$Q=\int{\rm d}\Pi_{b}\,\omega_{b}\!\left[\prod_{i=1}^{4}\!\int\frac{{\rm d}^{3}\vec{p}_{i}}{2\omega_{i}(2\pi)^{3}}\right]f_{1}f_{2}(1-f_{3})(1-f_{4})(2\pi)^{4}\delta^{4}(p_{1}+p_{2}-p_{3}-p_{4}-p_{b})\,\frac{\sum_{s}|\mathcal{M}_{\mathrm{SC}}|^{2}}{4},$ (A.3)
@@
${\rm d}^{3}\vec{p}_{i}=|\vec{p}_{i}|^{2}{\rm d}|\vec{p}_{i}|{\rm d}\Omega_{i}\simeq p_{F}m_{\psi}{\rm d}\omega_{i}{\rm d}\Omega_{i}=(m_{\psi}p_{F}T)\,{\rm d}y_{i}{\rm d}\Omega_{i},$ (A.4)
@@
$\begin{split}Q_{a}&=\frac{p_{F}^{4}}{2^{16}\pi^{10}}\!\left(\frac{256}{3}\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\right)A^{\mathrm{SC}}B^{\mathrm{SC}},\\[3.0pt] Q_{\mathrm{EDM/MDM}}&=\frac{p_{F}^{4}}{2^{16}\pi^{10}}\!\left(\frac{512}{3}\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{\mathrm{EDM/MDM}}^{\,2}}{m_{\pi}^{4}}\right)A^{\mathrm{SC}}B^{\mathrm{SC}},\end{split}$ (A.5)
@@
$\begin{split}A^{\mathrm{SC}}&=\!\left[\prod_{i=1}^{4}\!\int{\rm d}\Omega_{i}\right]\delta^{3}(\vec{p}_{1}+\vec{p}_{2}-\vec{p}_{3}-\vec{p}_{4})\!\left(\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}+\{\mathcal{M}_{kl}^{2}\}\right),\\[5.0pt] B^{\mathrm{SC}}&=T^{4}\!\!\int_{0}^{\infty}\!\!{\rm d}\omega_{b}\,\omega_{b}^{2}\!\left[\prod_{i=1}^{4}\!\int_{-\mu_{i}/T}^{\infty}\!\!{\rm d}y_{i}\right]\!\delta(\omega_{1}+\omega_{2}-\omega_{3}-\omega_{4}-\omega_{b})f_{1}f_{2}(1-f_{3})(1-f_{4}),\end{split}$ (A.6)
@@
$B^{\mathrm{SC}}$ $\approx T^{6}\!\!\int_{0}^{\infty}\!{\rm d}y_{b}\,y_{b}^{2}\!\left[\prod_{i=1}^{4}\!\int_{-\infty}^{\infty}\!{\rm d}y_{i}\,f_{i}\right]\delta\!\left(\sum_{i}y_{i}-y_{b}\right)=\frac{T^{6}}{6}\!\int_{0}^{\infty}\!{\rm d}y_{b}\,\frac{y_{b}^{3}(y_{b}^{2}+4\pi^{2})}{e^{y_{b}}-1}=\frac{62\pi^{6}T^{6}}{945}.$ (A.7)
@@
$F(\xi)$ $=4-\frac{1}{1+\xi^{2}}+5\xi\arctan(\xi)+\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}}\arctan\!\left(\frac{\xi^{2}}{\sqrt{1+2\xi^{2}}}\right)+\frac{\pi}{2}\!\left(\frac{\xi}{\sqrt{2\xi^{2}+1}}-5\right)\xi.$ (A.9)
@@ Appendix B Stellar Superradiance
$\begin{split}&\sum_{s}\left|\mathcal{M}^{a}_{\rm SS}\right|^{2}=\Bigg[\left\{\sum_{i=1}^{4}|\vec{v}_{i}|^{2}-4(\hat{l}\cdot\vec{v}_{1}+\hat{l}\cdot\vec{v}_{3})(\hat{l}\cdot\vec{v}_{2}+\hat{l}\cdot\vec{v}_{4})-4(\hat{k}\cdot\vec{v}_{1}+\hat{k}\cdot\vec{v}_{3})(\hat{k}\cdot\vec{v}_{2}+\hat{k}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{kl}^{2}\right\}\\ &+\sum_{i=1}^{4}|\vec{v}_{i}|^{2}\left(\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}\right)+\left\{2(\hat{k}\cdot\vec{v}_{1})(\hat{k}\cdot\vec{v}_{3})+2(\hat{k}\cdot\vec{v}_{2})(\hat{k}\cdot\vec{v}_{4})-(\vec{v}_{1}\cdot\vec{v}_{3})-(\vec{v}_{2}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{k}^{2}\right\}\\ &+\left\{2(\hat{l}\cdot\vec{v}_{2})(\hat{l}\cdot\vec{v}_{3})+2(\hat{l}\cdot\vec{v}_{1})(\hat{l}\cdot\vec{v}_{4})-(\vec{v}_{2}\cdot\vec{v}_{3})-(\vec{v}_{1}\cdot\vec{v}_{4})\right\}\left\{\mathcal{M}_{l}^{2}\right\}\Bigg]\,\frac{g_{\pi N}^{4}m_{\psi}^{2}g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.\end{split}$ (B.1)
@@
$\braket{\vec{k}\cdot\vec{q}_{b}}=\braket{\vec{l}\cdot\vec{q}_{b}}=\braket{\vec{p}_{1}\cdot\vec{q}_{b}}=0,\quad$ (B.2)
@@
$\left\langle\sum_{\text{s}}\left|\mathcal{M}^{a}_{\rm SS}\right|^{2}\right\rangle=\left[\frac{256}{3}\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}+128\frac{p_{F}^{2}}{m_{\psi}^{2}}\right]\left[\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right]\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{a}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.$ (B.3)
@@
$\varepsilon_{+1}^{\mu}(q)\varepsilon_{+1}^{*\nu}(q)\approx\frac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&i&0\\ 0&-i&1&0\\ 0&0&0&0\end{pmatrix}$, (B.4)
@@
$\left\langle k_{x}\,p_{1x}\right\rangle=-\frac{|\vec{k}|^{2}}{6},$ (B.5)
@@
$\left\langle\sum_{s}\big|\mathcal{M}^{\rm MDM}_{\rm SS}\big|^{2}\right\rangle=$ $\;\Bigg\{\frac{|\vec{q}_{b}|^{2}}{m_{b}^{2}}\bigg[\frac{4}{3}\{\mathcal{M}_{k}^{2}\}+\frac{4}{3}\{\mathcal{M}_{l}^{2}\}+\frac{8}{9}\{\mathcal{M}_{kl}^{2}\}\bigg]+\frac{2}{3}\frac{p_{F}^{2}}{m_{\psi}^{2}}\big[\{\mathcal{M}_{k}^{2}\}+\{\mathcal{M}_{l}^{2}\}\big]$ (B.6)
@@
$\left\langle\sum_{s}|\mathcal{M}^{\mathrm{EDM}}_{\rm SS}|^{2}\right\rangle=\frac{256}{3}\left(\left\{\mathcal{M}_{k}^{2}\right\}+\left\{\mathcal{M}_{l}^{2}\right\}+\left\{\mathcal{M}_{kl}^{2}\right\}\right)\frac{g_{\pi N}^{4}m_{\psi}^{2}\,g_{\rm EDM}^{2}}{m_{\pi}^{4}}\frac{m_{b}^{2}}{\Gamma_{\rm col}^{2}}.$ (B.7)
B.1 Phase Space Integral